A non-interest bearing promissory note has a face value of $950. Find the proceeds of this note if it is discounted 3½ years before its maturity date at 8% compounded quarterly


1
Expert's answer
2021-10-18T15:21:37-0400

The formula for calculating the value of the note:




"P=S(1-\frac{d}{m})^{m\cdot n},"

where "P" is the value of the note;

"S" - face value;

"d" - interest rate for the settlement period (day, month, year, ...);

"m" is number of charges per year.

"P=950(1-\frac{0.08}{4})^{4\cdot 3.5}=715.96" $

Discount of this note "S-P=950-715.96=234.04" $